A gallon of Del Monte White Vinegar is labeled 5%acidity. How many fluid ounces of pure acid must be added to get 6% acidity?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

A gallon of Del Monte White Vinegar is labeled 5%acidity. How many fluid ounces of pure acid must be added to get 6% acidity?
-----
Add g gallons of 100% acid
------
1*5 + g*100 = (1+g)*6
100g + 5 = 6g + 6
94g = 1
g = 1/94 gallons
Convert to fluid ounces (one of the dumbest units of volume on the planet)

Answer by stanbon(75887)

A gallon of Del Monte White Vinegar is labeled 5%acidity. How many fluid ounces of pure acid must be added to get 6% acidity?
------
Equation:
acid + acid = acid
1 gallon = 128 ounces
----------
0.05*128 + 1*x = 0.06(128+x)
-----
Multiply thru by 100 to get:
5*128 + 100x = 6*128 + 6x
94x = 128
x = 1.3617 ounces (amt. of pure acid to add)
==============
Cheers,
Stan H.
===============